# Cycloidal Quasi-Direct Drive Actuator Designs with Learning-based Torque Estimation for Legged Robotics

### v1

@@ I Introduction
Agile and dynamic locomotion in legged robots has been a longstanding challenge in robotics, requiring actuators that can deliver both dynamic torque control and high responsiveness, such as Direct Drive (DD) [1], [2] and Quasi-Direct Drive (QDD) [3] systems. However, integrating the gearbox into a confined space with conventional planetary, spur gear, and belt drive mechanisms is challenging without sacrificing gear load capacity since they are less resilient to significant impulse load, such as those experienced during a fall.
@@
Data-driven approaches, such as the Actuator Net [8] for modeling C-QDD, offer advantages since it is challenging to measure and account for manufacturing tolerances, to which cycloidal gears are particularly sensitive [9]. The significance of learning-based modeling lies in its ability to generalize across different operating conditions while maintaining high accuracy using only actuation history data [10].
@@
This paper presents a QDD actuator with a 10:1 cycloidal gearbox for legged robots. We also present a gated recurrent unit (GRU) based torque estimation framework to model nonlinearities induced by the cycloidal gear such as the torque ripple. The C-QDD actuator is benchmarked and the performance of the GRU torque estimation framework is evaluated experimentally.
@@
    Hardware verifications: We verify and analyze the C-QDD performances, backdrivability, and the torque estimation framework on hardware.
II Related Works
II-A Actuation Mechanism in Legged Robotics
Traditional robotic actuators are designed to be rigid, enabling precise position control [12], [13]. However, rigid and non-backdrivable actuators are not ideal for dynamic and agile-limbed systems [14], [15], as compliance and torque control are essential for legged robots [3]. Various types of actuators have been proposed to enable accurate force control and safe interaction with the environment in legged robots. Compliant actuators, such as series elastic actuators [16] and variable stiffness actuators [17], address these challenges by introducing a compliant element between the motor and the output. However, these actuators require extensive modeling, optimization, and complex control algorithms. Recently, QDD actuators have gained popularity in legged robots [18]. These proprioceptive actuators offer high torque fidelity and reduce reflected inertia, making them an ideal choice for legged robots. DD[19] and QDD [20] actuators have been foundational in achieving precise torque control and high responsiveness in legged robots. DD actuators are praised for their simplicity and torque transparency, which minimize mechanical impedance [3] and allow for accurate force control. In contrast, QDD actuators integrate gear reductions with direct drive principles, balancing torque amplification and control transparency [21]. Recent advancements in QDD technology have focused on optimizing gear ratios and reducing backlash, further enhancing the actuator’s performance in high-impact environments [22].
II-B Cycloidal Gear Mechanisms
Cycloidal gears have gained popularity as an alternative to planetary gears in legged robots due to their capacity to withstand large dynamic loads and frequent impacts [5], [23]. Additionally, cycloidal gearboxes are compact and typically exhibit less backlash compared to planetary gears [24]. However, despite these advantages, cycloidal gears also present several drawbacks stemming from the unique dynamics of the system [25]. These include torque ripple, backlash, and unintended vibrations caused by the eccentricity of the rotors, which can be compensated by using data-driven modelling.
II-C Actuator Modeling
Advanced actuator modeling and control are essential for managing the complexities of modern actuator designs, which introduce significant non-linearities and dynamic variations [26]. Traditional control methods often struggle with these challenges, but data-driven approaches, such as Actuator Networks, have proven effective in improving real-time control and prediction accuracy. By utilizing machine learning to model actuator dynamics based on offline data, Actuator Networks help bridge the sim-to-real gap in legged robots, as demonstrated by ETH Zurich’s ANYmal equipped with Series Elastic Actuators (SEA), enhancing both adaptability and performance [8]. Moreover, integrating RL with these actuator designs enable robots to optimize control policies for complex tasks autonomously. However, RL still faces challenges such as unsafe exploration and sim-to-real transfer difficulties [27], [28]. Actuator Networks address these challenges by improving the accuracy of actuator models, thereby aiding in the transferability of learned policies from simulation to real-world environments [29].
III Cycloidal Gear Quasi-Direct Drive Mechanisms
@@
The use of cycloidal gears in low gear ratio and high backdrivability actuators introduces unique challenges to the design of cycloidal mechanisms. In this section, we will discuss the design principles behind C-QDD actuators, focusing on how these parameters are optimized to achieve the desired performance characteristics.
@@ III-A Sizing of Cycloidal and Planetary Gears
Here, we compare the load capacity of our cycloidal gears with that of traditional planetary gears, which are commonly used in proprioceptive actuators [5]. The gearbox is integrated into the stator’s interior to maintain a compact design, limiting the maximum allowable pitch diameter for the cycloidal gear. A compound planetary gear can achieve a gear ratio of $10$ while remaining compact enough to fit within the stator. One such configuration includes a sun gear with $12$ teeth, planet gears with $36$ and $22$ teeth, and a ring gear with $70$ teeth.
@@ III-B Cycloid profiles and eccentricity
where $Z_{np}$ represents the number of outer pins, and $Z_{nt}$ denotes the number of teeth on the cycloid gear. The number of disks, $N_{Z}$ required is denoted as
@@
$k$ $\in\mathbb{Z},\quad k\geq 2$ (4)
Here, $k$ is a positive integer, and $N_{\text{disks}}$ is the number of counterbalancing disks.
Two disks are sufficient to balance the inertia statically. However, to ensure dynamic stability, three disks are necessary to compensate for dynamic imbalances that may arise due to material deformation and manufacturing inaccuracy [9]. In the C-QDD design, two counter disks are employed as the transmission ratio, $Z_{R}=10$, necessitates the use of an even number of disks and significantly increases reflected inertia, impacting the backdrivability of the actuator. The dynamic imbalance of the cycloid disk contributes to oscillations and torque ripples, which are compensated for through the torque estimation framework in Section V-C.
@@ III-C Non-Pinwheel Cycloid Ring
In the C-QDD, a non-pinwheel outer cycloid ring is employed, instead of the conventional pinwheel designs [30], [31]. Both the cycloid disks and the outer rings are manufactured from $4140$ alloy steel, known for their superior rigidity, wear resistance, and dimensional stability compared to the $7075$ aluminum housing. Pinwheel designs traditionally utilize dowel pins embedded in the housing, which can present manufacturing challenges. Drilling pinholes directly into the housing necessitates tight tolerance control. C-QDD resolves these challenges by utilizing independent solid rings, which may accommodate better manufacturing tolerances among the cycloid disks, eccentric cams, and other components. The cycloid disks and rings are machined from tight tolerance ground flat stock using a wire EDM at ±$\pm 2\,\mu\text{m}$ tolerance.
Figure 4: Overview of the Torque Estimation Framework for torque estimation.
IV C-QDD System Dynamics and Torque Estimation
@@ IV-A Torque Estimation Framework
The proposed torque estimation framework utilizes a Gated Recurrent Unit (GRU) [32] based architecture to model the actuator’s dynamic behavior through the joint states. Over the time history horizon, $L$, two different sets of input vectors are defined as:
  • PV-GRU: $\mathbf{X}_{v_{t}}\in\mathbb{R}^{2L}$, where $\mathbf{X}_{v_{t}}=[\mathbf{q_{e}},\dot{\mathbf{q}}]^{\top}$
  • PVA-GRU: $\mathbf{X}_{at}\in\mathbb{R}^{3L}$, where $\mathbf{X}_{a_{t}}=[\mathbf{q_{e}},\dot{\mathbf{q}},\ddot{\mathbf{q}}]^{\top}$
@@
The GRU architecture was selected due to its capability to capture temporal dependencies efficiently, even across sequences of varying lengths. By maintaining an internal memory, the GRU processes historical state information $x_{t}$ along with the hidden state $h_{t-1}$ to iteratively update its predictions, thus enabling the network to learn long-term dependencies that are critical for accurately predicting high-frequency torque ripple.
IV-B System Identification and Training
The network is trained to minimize the error between the predicted torque and the actual measured torque, $\tau_{pred}$. Supervised learning is employed using historical data collected under diverse operating conditions of the C-QDD actuator, capturing a wide range of dynamic behaviors.
The training pipeline uses the Adam optimizer with an initial learning rate of $0.0001$, coupled with a One Cycle Learning Rate (LR) scheduler for efficient convergence [33], [34]. This approach dynamically adjusts the learning rate to prevent overfitting while fine-tuning model performance during the final training stages.
Key hyperparameters were tuned to optimize performance:
  • History size ($L=30$)—providing sufficient temporal information to capture high-frequency dynamics.
  • Number of GRU layers ($K=4$)—striking a balance between model complexity and computational efficiency.
  • Batch size = 64.
The parameter $L$ was determined by considering the dataset’s sampling rate and the observed torque ripple frequency range. This ensures that the network receives enough historical data to capture the actuator’s intricate dynamics. Similarly, $K$ was adjusted to enable the network to model higher-order features while maintaining computational efficiency effectively, crucial for real-time applications.
@@ V-A C-QDD Mechanical and Electrical Performance
C-QDD’s continuous and peak torque values were measured using a high-impedance set-up and evaluated at $37.5$ Nm and $89.9$ Nm, respectively. The peak torque is significantly lower than the expected peak torque of $120$ Nm based on the BLDC due to current limits on the C-QDD’s driver board. Table. V-A compares C-QDD to some of the actuators intended for legged robots with similar features. A low amplitude torque tracking bandwidth was also experimentally obtained and evaluated at 34.3 Hz. This was done through the high-impedance setup and by commanding a torque chirp signal to the actuator with a magnitude of 5 Nm.
TABLE I: C-QDD perfromance metrics
@@ V-B1 Backlash
The backlash of C-QDD was measured based on the method discussed in [36]. For this test, the C-QDD chassis was fixed to the bed with no load attached to its output. The actuator is first preloaded in the opposite direction, then the BLDC motor is gradually rotated until an output shaft motion is observed. A dial indicator with an accuracy of $0.0127$ mm was used to identify when the output shaft moved. To account for inaccuracies in the manufacturing and assembly processes, this test was performed at six different locations of the output shaft, preloaded in both directions. The average backlash was calculated at $7.0$ arcminutes with an uncertainty of ±$\pm 1.3$ arcminutes. The backlash is comparable to other actuators designed for similar purposes, such as the one in [22].
@@ V-B2 Static and Dynamic Backdriving Torque
The torque needed to back-drive the C-QDD from the output was measured through the test setup shown in Fig. 5. The servo motor was used to back-drive the output of the C-QDD with a torque sensor rigidly fixed concentrically between the outputs of the actuator and servo. The static and dynamic backdriving torques—required to initiate and maintain motion, respectively—were both evaluated. The static and dynamic backdriving torques were observed at $1.99$ Nm and $1.36$ Nm, respectively. When compared to other QDD actuators, such as [22] and [37], which have backdriving torques of $0.37$ Nm and $0.97$ Nm, respectively, the backdriving force is higher, but comparable considering the C-QDD incorporates a higher gear-ratio and a considerately higher torque capability.
Figure 6: The estimated torque and the ground truth torque
@@ V-C Torque Estimation and Actuator Network Performance
History Size 30 30 24 3
*Cycle time refers to the execution time on the fastest available device: GPU for the GRU models and CPU for the MLP models.
To evaluate the Actuator Network’s performance for torque estimation, $\mathcal{X}=\{q,\dot{q},\ddot{q},\tau_{q}\}$ were collected on a single inverted pendulum setup using our C-QDD. The training data was generated using sinusoidal joint angle profiles with varying frequencies $f_{q}\in(0.3,1.5)$ Hz, amplitude $\mathcal{A}_{q}\in(0.5,1.5)$ rad, and initial position $q_{t=0}\in(0,1)$ rad. The pendulum mass, $m_{p}\in\{1.14,2.28\}$ kg was attached at varying locations on the pendulum bar, $r_{p}\in(0.25,0.6)$ m, which represents various loads and inertia of the pendulum arm. The pendulum is allowed to make contact with a compliant wall, simulating the hybrid dynamics characteristics of legged robot motion.
V-C1 Actuator Net Performance Comparisions
@@
As a baseline, a multilayer perceptron (MLP) used in [8] was trained on our dataset, and a tuned version of the MLP was also trained for further comparison. The tuned MLP and baseline MLP had an RMSE of $1.711$ Nm and $2.457$ Nm, respectively, demonstrating comparable results to the original implementation in [8], which is applied to a serial elastic actuator instead of a cycloid gear and operates under smaller torque ranges. Fig. 6 shows a visualization where all models agree with the ground truth for large-scale torque estimation. However, in Fig. 6b, the tuned MLP could not accurately capture small torque variation features, such as sinusoidal torque ripple with $0.6$ Nm amplitude and $20$ Hz frequency, showing a phase shift behavior of almost $180\degree$.
On the other hand, the PVA-GRU model effectively leveraged the higher-order acceleration data to predict small amplitude and high-frequency variations accurately. As shown in Fig. 6b, the PVA-GRU estimated the torque ripple amplitude to be on average within $0.23$ Nm, with a frequency of $17.61$ Hz, compared to the ground truth value of $17.24$ Hz, and a phase shift of less than $40\degree$. Furthermore, the model demonstrated a significant improvement of $43.54\%$ in RMSE and $68.21\%$ in variance compared to the tuned MLP. Accurately capturing such torque nonlinearities is crucial for applications like RL simulations, where minor discrepancies in torque estimation can impact performance [7].

### v2

@@ I Introduction
Agile and dynamic locomotion in legged robots has been a longstanding challenge in robotics, requiring actuators that can deliver both dynamic torque control and high responsiveness, such as Direct Drive (DD) [1], [2] and Quasi-Direct Drive (QDD) [3] systems. However, integrating the gearbox into a confined space with conventional planetary, spur gear, and belt drive mechanisms is challenging without sacrificing gear load capacity since they are less resilient to significant impulse loads, such as those experienced during a fall.
@@
Data-driven approaches for torque and force estimation, such as the Actuator Net [8], [9], offer advantages since it is challenging to measure and account for manufacturing tolerances, to which cycloidal gears are particularly sensitive [10]. The significance of learning-based modeling lies in its ability to generalize across different operating conditions while maintaining high accuracy using only actuation history data [11].
@@
This paper presents a QDD actuator with a 10:1 cycloidal gearbox for legged robots, shown in Fig. 1. We also propose a torque estimation framework using gated recurrent units (GRU) [12] to model nonlinearities induced by the cycloidal gear, such as torque ripple. The C-QDD actuator is benchmarked and the performance of the torque estimation framework is evaluated experimentally.
@@
    Hardware verifications: We evaluate the performance of the C-QDD and torque estimation framework through hardware experiments.
@@
II Related Works
II-A Actuation Mechanism in Legged Robotics
Traditional robotic actuators are designed to be rigid and for position control [14], [15]. However, rigid and non-backdrivable actuators are not ideal for dynamic and agile motion [16], [17], as compliance and torque control are essential in such scenarios [3]. Compliant actuators, such as series elastic [18] and variable stiffness actuators [19], address these challenges with a compliant spring element between the motor and the output. However, they require extensive modeling, optimization, and complex control algorithms. Proprioceptive actuators, such as DD [20] and QDD [21] actuators have been foundational in achieving high torque control fidelity and high responsiveness in legged robots [22]. DD actuators are praised for their simplicity and torque transparency, which minimize mechanical impedance [3] and lower reflected inertia. In contrast, QDD actuators integrate gear reductions, while balancing torque amplification and maintaining the benefit of DD [23]. Recent advancements in QDD technology have focused on optimizing gear ratios and reducing backlash, further enhancing the actuator’s performance in high-impact environments [24].
II-B Cycloidal Gear Mechanisms
Cycloidal gears have gained popularity as an alternative to planetary gears in legged robots due to their capacity to withstand large dynamic loads and frequent impacts [5], [25]. Cycloidal gearboxes are compact and typically exhibit less backlash compared to planetary gears [26]. However, despite these advantages, cycloidal gears also present several drawbacks stemming from the unique dynamics of the system [27]. These include torque ripple, backlash, and unintended vibrations caused by the eccentricity of the rotors.
II-C Actuator Modeling
Actuator modeling and control are essential since the actuator dynamics inevitably contain non-linearity and dynamic variations [28]. Actuator system identification is becoming increasingly important for addressing the sim-to-real gap in RL [29], [30]. Traditional methods often struggle with these challenges, while data-driven approaches, such as actuator networks [8], have proven effective in improving the motor sim-to-real gap and torque prediction accuracy [31].
Our C-QDD aims to benefit from the cycloid gear and quasi-direct drive mechanisms while compensating for the cycloid’s complex dynamics by using a data-driven approach.
Figure 4: Overview of the torque estimation framework architecture, leveraging temporal
actuator state information to generate precise torque predictions.
III Cycloidal Gear Quasi-Direct Drive Mechanisms
The use of cycloidal gears in low gear ratio and highly backdrivable actuators introduce unique challenges to the design of cycloidal mechanisms. In this section, we discuss the design principles behind C-QDD actuators displayed in Fig. 2 and Fig. 3, focusing on how these parameters are optimized to achieve the desired performance characteristics.
@@ III-A Sizing of Cycloidal and Planetary Gears
Here, we compare the load capacity of our cycloidal gears to traditional planetary gears, which are common in proprioceptive actuators [5]. The gearbox is integrated into the stator’s interior to maintain a compact design, limiting the maximum allowable pitch diameter for the cycloidal gear. A compound planetary gear can achieve a gear ratio of $10$ while remaining compact enough to fit within the stator. One such configuration includes a sun gear with $12$ teeth, planet gears with $36$ and $22$ teeth, and a ring gear with $70$ teeth.
@@ III-B Cycloid profiles and eccentricity
where $Z_{np}$ represents the number of outer pins, and $Z_{nt}$ denotes the number of teeth on the cycloid gear. The number of disks, $N_{Z}$ required is denoted as:
@@
Here, $k\in\mathbb{Z},\quad k\geq 2$, and $N_{\text{disks}}$ is the number of counterbalancing disks.
Two disks are sufficient to balance the inertia statically. However, to ensure dynamic stability, three disks are necessary to compensate for dynamic imbalances that may arise due to material deformation and manufacturing inaccuracy [10]. In the C-QDD design, two counter disks are used because the transmission ratio $Z_{R}=10$ requires an even number of disks. However, this also significantly increases reflected inertia, which impacts the actuator’s backdrivability. The dynamic imbalance of the cycloid disk contributes to oscillations and torque ripples, which are compensated for through the torque estimation framework in Section V-C.
@@ III-C Non-Pinwheel Cycloid Ring
In the C-QDD, a non-pinwheel outer cycloid ring is employed, instead of the conventional pinwheel design [32], [33]. Both the cycloid disks and the outer rings are manufactured from $4140$ alloy steel, known for their superior rigidity, wear resistance, and dimensional stability compared to the $7075$ aluminum housing. Pinwheel designs traditionally utilize dowel pins embedded in the housing, which can present manufacturing challenges. Drilling pinholes directly into the housing necessitates tight tolerance control. C-QDD resolves these challenges by utilizing independent solid rings, which may accommodate better manufacturing tolerances among the cycloid disks, eccentric cams, and other components. The cycloid disks and rings are machined from tight tolerance ground flat stock using a wire EDM at ±$\pm 2\,\mu\text{m}$ tolerance.
IV C-QDD Dynamics and Torque Estimation
@@ IV-A Torque Estimation Framework
The proposed torque estimation framework utilizes a GRU-based architecture to model the actuator’s dynamic behavior through the joint states. Over the time history horizon, $L$, two different sets of input vectors are defined as:
  • PV-GRU: $\mathbf{X}_{v_{t}}\in\mathbb{R}^{2\times L}$, where $\mathbf{X}_{v_{t}}=[\mathbf{q_{e}},\dot{\mathbf{q}}]^{\top}$
  • PVA-GRU: $\mathbf{X}_{at}\in\mathbb{R}^{3\times L}$, where $\mathbf{X}_{a_{t}}=[\mathbf{q_{e}},\dot{\mathbf{q}},\ddot{\mathbf{q}}]^{\top}$
@@
The GRU architecture was selected because of its capability to capture temporal dependencies efficiently, even across sequences of varying lengths. By maintaining an internal memory, the GRU processes historical state information $x_{t}$ along with the hidden state $h_{t-1}$ to iteratively update its predictions, thus enabling the network to learn long-term dependencies that are critical for accurately predicting high-frequency torque ripple.
IV-B Model Parameters and Training
The network is trained to minimize the error between the predicted torque, $\tau_{pred}$, and the actual measured torque. Supervised learning is employed using historical data collected under diverse operating conditions of the C-QDD actuator, capturing a wide range of dynamic behaviors. The training pipeline uses the Adam optimizer with an initial learning rate (LR) of $1.0e^{-4}$, coupled with a One Cycle Learning Rate scheduler for efficient convergence [34], [35]. This approach dynamically adjusts the LR to prevent overfitting while fine-tuning model performance during the final training stages.
For our C-QDD, a $4$-layer stacked GRU with history size $30$ and batch size $64$ are selected. The history size was determined by considering the dataset’s sampling rate—200Hz—and the observed torque ripple frequency range. This ensures that the network receives enough historical data to capture the actuator’s intricate dynamics. Similarly, the GRU layers were tuned to introduce sufficient complexity in the network to model higher-order features while maintaining computational efficiency, crucial for real-time applications.
@@ V Hardware Verification and Analysis
Figure 6: Torque Estimation Framework data collection and performance testing setup.
@@ V-A C-QDD Mechanical and Electrical Performance
C-QDD’s continuous and peak torque values were measured using a high-impedance set-up and evaluated at $37.5$ Nm and $89.9$ Nm, respectively. The peak torque is lower than the expected peak torque of $120$ Nm based on the BLDC due to current limits on the C-QDD’s CubeMars Driver Board-V2.1. Table. V-A compares C-QDD to some actuators intended for legged robots with similar features. A low amplitude torque tracking bandwidth was also experimentally obtained and evaluated at $34.3$ Hz. The torque control bandwidth was measured using the high-impedance setup by applying a $5$ Nm torque chirp signal to the actuator. The actuator output power was measured with a low-impedance dynamic torque transducer and encoder values. The peak electrical to mechanical output efficiency reached $82.3$%.
TABLE I: C-QDD performance metrics
@@ V-B1 Backlash
The backlash of the C-QDD was measured based on the method discussed in [37]. For this test, the C-QDD chassis was fixed to the bed with no load attached to its output. The actuator is first preloaded in the opposite direction, and then the BLDC motor is gradually rotated until an output shaft motion is observed. A dial indicator with an accuracy of $0.0127$ mm was used to identify when the output shaft moved. To account for inaccuracies in the manufacturing and assembly processes, this test was performed at six different locations of the output shaft, preloaded in both directions. The average backlash was calculated at $7.0$ arcminutes with an uncertainty of ±$\pm 1.3$ arcminutes. The backlash is comparable to other actuators designed for similar purposes, such as the one in [24].
@@ V-B2 Static and Dynamic Backdriving Torque
The torque needed to back-drive the C-QDD from the output was measured through the test setup shown in Fig. 5. The servo motor was used to back-drive the output of the C-QDD with a torque sensor rigidly fixed concentrically between the outputs of the actuator and servo. The static and dynamic backdriving torques—required to initiate and maintain motion, respectively—were both evaluated. The static and dynamic backdriving torques were observed at $1.99$ Nm and $1.36$ Nm, respectively. When compared to other QDD actuators, such as [24] and [38], which have backdriving torques of $0.37$ Nm and $0.97$ Nm, the backdriving force is higher, but comparable considering the C-QDD incorporates a higher gear-ratio and has a considerately higher torque capability.
Figure 7: The graphs show the estimated torque for all the models compared to the ground truth torque during live inference, along with their variances.
@@ V-C Torque Estimation and Actuator Network Performance
To evaluate the Actuator Network’s performance for torque estimation, $\mathcal{X}=\{q,\dot{q},\ddot{q},\tau_{q}\}$ were collected on a single inverted pendulum setup using our C-QDD shown in Fig. 6. The training data was generated using sinusoidal joint angle profiles with varying frequencies $f_{q}\in[0.3,1.5]$ Hz, amplitude $\mathcal{A}_{q}\in[0.5,1.5]$ rad, and initial position $q_{t=0}\in[0,1]$ rad. The pendulum mass, $m_{p}\in\{1.14,2.28\}$ kg was attached at varying locations on the pendulum bar, $r_{p}\in[0.25,0.6]$ m, which represents various loads and inertia of the pendulum arm. The pendulum is allowed to make contact with a compliant wall, simulating the hybrid dynamics characteristics of legged robotic motion.
@@
* Cycle time refers to the execution time on the fastest device; device is CPU for all models; all models use a hidden size of 32.
V-C1 Actuator Net Performance Comparisons
@@
As a baseline, a multilayer perceptron (MLP) used in [8] was trained on our dataset, and a tuned version of the MLP was also trained for further comparison. The tuned MLP and baseline MLP had an RMSE of $1.71$ Nm and $2.46$ Nm, respectively, demonstrating comparable results to the original implementation in [8], which is applied to a serial elastic actuator instead of a cycloidal gear actuator and operates under smaller torque ranges. Fig. 7a shows a visualization where all models agree with the ground truth for large-scale torque estimation. However, in Fig. 7b, the tuned MLP could not accurately capture small torque variation features, such as sinusoidal torque ripple with $0.6$ Nm amplitude and $20$ Hz frequency, showing a phase shift behavior of almost $180\degree$.
On the other hand, the PVA-GRU model effectively leverags the higher-order acceleration data to predict small amplitude and high-frequency variations accurately. As shown in Fig. 7b, the PVA-GRU estimated the torque ripple amplitude with an average of $0.23$ Nm error from the ground truth, with a frequency of $17.6$ Hz, compared to the ground truth of $17.2$ Hz, and a phase shift of less than $40\degree$. Furthermore, the model demonstrated a significant improvement of $43.5\%$ in RMSE and $68.2\%$ in variance compared to the tuned MLP. Accurately capturing such torque nonlinearities is crucial for applications like RL simulations, where minor discrepancies in torque estimation can impact performance [7].